Bill is 5 times older than Keira, Bill is 5 times older than Phil, Oscar is twice as old as Bill,
and the total age of all of them at 81. How old is Phil?

Found 4 solutions by ikleyn, greenestamps, ankor@dixie-net.com, josgarithmetic:
Answer by ikleyn(53756)

.
Bill is 5 times older than Keira, Bill is 5 times older than Phil, Oscar is twice as old as Bill,
and the total age of all of them at 81. How old is Phil?
~~~~~~~~~~~~~~~~~~~


In English,  the forms like this   "Bill is 5 times older than Keira"   are prohibited in  Math problems formulations.


In this part,  English is very different from many other languages.


            (In this part,  English is  RADICALLY  different from many other languages,
            and the newcomers,  as a rule,  don't know and even don't suspect about it  ( ! ) )


Therefore,  when you translate from your native language,  be careful -- you should know standard English patterns
for the subject to which you create/translate your problem.



Answer by greenestamps(13332)



ANSWER: NO SOLUTION!

The statement of the problem is faulty. Age problems need to have answers that are whole numbers; this one does not -- regardless of how it is interpreted.

The statement of a math problem should never say that one person is "5 times OLDER THAN" another person.

In sloppy everyday English usage, "5 times OLDER THAN" and "5 times AS OLD AS" are used to mean the same thing, but they do not. If the first person's age is x, then the age of a person 5 times AS OLD AS the first is clearly 5 times x, or 5x. But if the age of the second person is 5 times OLDER THAN the first, then that person's age is x, PLUS 5 MORE TIMES x, which is x+5x=6x.

This problem is faulty in that the answers are not whole numbers with either interpretation of the given information.

(1) If we incorrectly use 5 times older than x to mean 5x, then
Phil = x
Bill = 5x
Keira = x
Oscar = 10x

The sum of their ages is 81:
17x=81
x = 81/17 not a whole number

(2) if we correctly use 5 times older than x to mean 6x, then
Phil = x
Bill = 6x
Keira = x
Oscar = 12x

20x=81
x = 81/20 not a whole number


Answer by ankor@dixie-net.com(22740)

Bill is 5 times older than Keira,
b = 5k
Bill is 5 times older than Phil,
b = 5p
Oscar is twice as old as Bill,
o = 2b
and the total age of all of them at 81.
b + k + p + o = 81
Something wrong here, as written, this will not produce an integer solution which is generally required in problems like this.
Change a couple words in this problem, it works then
"
Bill is "3 years" older than Keira,
b = k + 3
k = b - 3
Bill is 5 times older than Phil,
b = 5p
divide both sides by 5
p = .2b
Oscar is twice as old as Bill,
o = 2b
and the total age of all of them at 81.
b + k + p + o = 81
replace: k with (b-3), replace p with .2b and o with 2b
b + (b-3) + .2b + 2b = 81
4.2b = 81 + 3
b = 84
b = 84/4.2
b = 20 yrs old is bill
then
k = 20-3 = 17 yrs is Keira
and
p = .2*20 = 4 yrs is Phil
and
o = 2 * 20 = 40 yrs is Oscar
:
Check: 20 + 17 + 4 + 40 = 81
:
Anyway this is the method.


Answer by josgarithmetic(39797)

Strange description.
If "Bill is 5 times older than Keira, Bill is 5 times older than Phil,", then Keira and Phil are the same age.
Bill        k+5k=p+5p
Keira       k
Phil        p
Oscar      2(k+5k)=2(p+5p)
TOTAL       81

That table can be simplified because the simple addition and k=p.
Bill        6p
Keira       p
Phil        p
Oscar       12p
TOTAL       81

Question asked essentially for p.